1. Use the first principal to differentiate the following with respect to x

f(x) = 3x2- 4x + 7


1
Expert's answer
2021-05-11T07:02:02-0400

Solution. Using differentiation from first principles


"\frac{dy}{dx}=\lim\limits_{h\to0}\frac{f(x+h)-f(x)}{h}"

According to the condition of the problem f(x)=3x2-4x+7. Therefore


"f(x+h)=3(x+h)^2-4(x+h)+7=3x^2+6hx+3h^2-4x-4h+7."

Hence


"f(x+h)-f(x)=3x^2+6hx+3h^2-4x-4h+7-3x^2+4h-7""f(x+h)-f(x)=6hx+3h^2-4h"

As result get


"\frac{dy}{dx}=\lim\limits_{h\to0}\frac{6hx+3h^2-4h}{h}=\lim\limits_{h\to0}(6x+3h-4)=6x-4"

Answer.

"\frac{dy}{dx}=6x-4"